An electric power station that operates at 30 kV and uses a 20:1 step-up ideal transformer is producing 330 MW (Mega-Watt) of power that is to be sent to a big city which is located 300 km away with only 2.5% loss. Each of the two wires are made of copper (resistivity = 1.68 × 10 − 8 Ω.m). What is the diameter of the wires?

Please show detailed work!
Answer #1

V = 30 kV = 20 x 30 x 10^3 V = 6 x 10^5 V

VI = 330 MV

I = (330 x 10^6) / ( 6 x 10^5)

I = 550 A

2.5% loss = I^2R = 2.5/100 x 330 MW

550^2 x R = 8.25 x 10^6

R =  8.25 x 10^6 / 550^2

R = 27.27 ohms

d = (1.68 x 10^-8 x 300 x 10^3) / 27.27

d = 1.84 x 10^4 m